please help to solve for x. 9^x=1/3
Answer by stanbon(48569)

please help to solve for x. 9^x=1/3
-------------
Write both sides to the base 3.
--------
(3^2)^x = 3^-1
3^(2x) = 3^-1
2x = -1
x = -1/2
============
Cheers,
Stan H.